a man standing on railway platform listening to the whistle of an engine tat passes the man at constant speed at time "   t''   ".


Hoow does the frequency as heard by man changes wid time.


Sketch the graph of "f vs t".

Grade:12

1 Answers

vikas askiitian expert
13 years ago

frequency heard by listener from t=0 (when train was approaching and was far apart ) to t=T (when train passes the listener) is constant & given by

f = fo(V/V-Vt)                      (Vt = speed of train, V speed of sound)

when train has passed the listener then frequency again becomes constant but now it is f1

f1 = fo(V/V+Vt)

now plot this graph by using these two eq